OpenStudy (anonymous):

Solve: -5(x - 2) = 15 Answer -5 -3.4 1 -1

jhonyy9 (jhonyy9):

-5x+10=15 -5x=5 x=5/(-5) x= -1

OpenStudy (anonymous):

-1

OpenStudy (ars22):

-1

OpenStudy (anonymous):

-5x+10 =15 -5x=5 x=5/-5 x= -1